# positioning

### v1

@@ Abstract.
We investigate and compare applications of the Zilber–Pink conjecture and dynamical methods to rigidity problems for arithmetic real and complex hyperbolic lattices. Along the way we obtain new general results about reconstructing a variation of Hodge structures from its typical Hodge locus that may be of independent interest. Applications to Siu’s immersion problem are also discussed.
@@ 1. Introduction and Motivation
A cornerstone in the representation theory of lattices in Lie groups is Margulis’s superrigidity theorem [28, p. 2]. Indeed, the implications of superrigidity are profound, including arithmeticity of irreducible lattices in all noncompact semisimple Lie groups except $\operatorname{\operatorname{PO}}(1,n)$ and $\operatorname{\operatorname{PU}}(1,n)$, i.e., for all but real and complex hyperbolic lattices. Specifically, Margulis proved superrigidity of irreducible lattices in semisimple Lie groups of real rank at least two. Using the theory of harmonic maps, this was later extended to the rank one groups $\mathrm{Sp}(1,n)$ and $\mathrm{F}_{4}^{(-20)}$ by Corlette [13] and Gromov–Schoen [21]. While the harmonic maps method was extended to give a new proof in higher rank [29], it is open (and quite relevant in context of the blend of techniques used in this paper) whether the dynamical methods of Margulis can give an independent proof in the rank one cases.
To be precise, let $\Gamma$ be a lattice in a Lie group $G$. A representation $\rho$ from $\Gamma$ to a topological group $H$ is said to extend if there is a continuous homomorphism $\widetilde{\rho}:G\to H$ so that $\rho$ is the restriction of $\widetilde{\rho}$ to $\Gamma$ under its lattice embedding in $G$. Then $\Gamma$ is superrigid if every unbounded, Zariski dense representation of $\Gamma$ to a connected, adjoint simple algebraic group $H$ over a local field $k$ of characteristic zero extends. It is known that there are lattices in $\operatorname{\operatorname{PO}}(1,n)$ and $\operatorname{\operatorname{PU}}(1,n)$ that are not superrigid; [22, Sec. 5] for real hyperbolic examples and [31, Sec. 22] or [24, Sec. 1] for some results in the complex hyperbolic setting. The study of representations of real and complex hyperbolic lattices is very difficult and rich in open questions, and the extent to which superrigidity fails has yet to be made precise. For example, one motivation for this paper is the first part of a question asked by David Fisher [19, Qtn. 3.15]:
@@
In particular, this paper builds on the techniques from [1, 2, 8] to explore geometric conditions on a representation that allow one to use tools from dynamics and/or Hodge theory to prove new rigidity phenomena. There is some history of success along these lines. For just one family of examples, some classes of representations studied in the literature have maximal geometric behavior, where two distinct instances are:
@@ 1.1. Siu’s immersion problem
An original motivation for studying the more general rigidity questions considered in this paper is Siu’s immersion problem [38, Prob. (b), p. 182]. Let $\mathbb{B}^{n}$ denote the unit ball in $\mathbb{C}^{n}$ with its Bergman metric. If $\Gamma$ is a lattice in the holomorphic isometry group $\operatorname{\operatorname{PU}}(1,n)$ of $\mathbb{B}^{n}$, then $S_{\Gamma}$ will denote the ball quotient $\Gamma\backslash\mathbb{B}^{n}$.
@@
To our knowledge the only previous result is by Cao and Mok [12], which answered 1.2 in the affirmative for $f:\Gamma\backslash\mathbb{B}^{n}\to\Lambda\backslash\mathbb{B}^{m}$ when $m<2n$. See [2, Thm. 1.7] and the surrounding references for more on Siu’s analogous submersive problem. The following is our main contribution toward a positive answer to Siu’s immersion problem.
@@
Let $f:S_{\Gamma_{1}}\to S_{\Gamma_{2}}$ be an immersive holomorphic map between arithmetic ball quotients of unequal dimension. Suppose that $S_{\Gamma_{1}}$ contains a proper totally geodesic subvariety of positive dimension. Then the following are equivalent:
@@
The unequal dimension hypothesis is simply to rule out coverings. The positive-dimensional hypothesis on the subvarieties of $S_{\Gamma_{1}}$ is to rule out the trivial counterexample of points, but note that $W_{\ell}$ could be a point. We offer two proofs of 1.3, one using homogeneous dynamics which gives a slightly weaker statement in Section 3, and the second using Hodge theory and functional transcendence in Section 4, as an application of 1.15 below.
The arithmetic ball quotients with totally geodesic subvarieties of every possible codimension are those of simplest type, which are constructed using hermitian forms over number fields. In this case, 1.3 reduces 1.2 to checking a single dimension.
@@
There are various similar generalizations that follow from 1.14. For example, a period map $\psi:S_{\Gamma_{1}}\to\Gamma\backslash D$ such that the $f^{-1}(Y_{i})$ are totally geodesic subvarieties of $S_{\Gamma_{1}}$ for a generic sequence of Mumford–Tate subdomains $Y_{i}\subset\Gamma\backslash D$ is an isomorphism.
@@ 1.2. Rich representations and the main theorem
Let $\Gamma\subset G$ be a torsion-free arithmetic lattice, where $G$ is $\operatorname{\operatorname{SU}}(1,n)$ or $\operatorname{\operatorname{SO}}(1,n)$, and let $X$ denote the symmetric space associated with $G$. Then $S_{\Gamma}$ will denote the manifold $\Gamma\backslash X$, which is a smooth manifold that is also a smooth variety in the complex hyperbolic case [^1] Since both contexts are relevant to this paper, we use the term submanifold when considering $S_{\Gamma}$ as a differentiable manifold and subvariety when $S_{\Gamma}$ is complex hyperbolic and the submanifold is also a subvariety for the natural algebraic structure.. We assume that $S_{\Gamma}$ contains one, and therefore infinitely many, proper totally geodesic submanifolds [^2] This claim follows easily from the fact that the associated algebraic group commensurates the lattice. In the other direction, if $S_{\Gamma}$ contains infinitely many maximal totally geodesic submanifolds of dimension at least $2$, then $\Gamma$ is arithmetic; see [1, 2, 8]. of real dimension at least $2$ (see Section 1.4 for the case of closed geodesics). When this is the case, we say that $S_{\Gamma}$ (or $\Gamma$) is (geodesically) rich. Let $\{S_{\Gamma_{i}}\}_{i\in\mathbb{N}}$ be a sequence of (immersed) maximal totally geodesic submanifolds, where maximal means that the only totally geodesic submanifold strictly containing $S_{\Gamma_{i}}$ is $S_{\Gamma}$ itself.
The group $\Gamma_{i}$ is the stabilizer in $\Gamma$ of a totally geodesic subspace $Y_{i}$ of $X$, and the restriction of $\Gamma_{i}$ to $Y_{i}$ induces a surjective homomorphism from $\Gamma_{i}$ onto the fundamental group of $S_{\Gamma_{i}}$. The kernel of this homomorphism encodes the action of $\Gamma_{i}$ on the normal bundle to $Y_{i}$ in $X$. The main purpose of this paper is to introduce the notion of (geodesically) rich representations, which are morally the ones that “preserve” a significant amount of the rich structure of $\Gamma$:
@@
Let $k$ be a local field and ${\mathbf{H}}$ be a semisimple $k$-algebraic group without compact factors. Suppose that $\Gamma$ is geodesically rich and that $\rho:\Gamma\to{\mathbf{H}}(k)$ is a representation with unbounded, Zariski dense image. Define ${\mathbf{H}}_{i}$ by
(1.2.1) $\mathbf{H}_{i}\coloneqq\overline{\rho(\Gamma_{i})}\subset{\mathbf{H}}$
for each $i$, where closure is with respect to the Zariski topology. Then $\rho$ is (geodesically) rich if there are infinitely many $i$ so that $\dim{\mathbf{H}}_{i}<\dim{\mathbf{H}}$.
@@
Let $\Gamma\subset G=\operatorname{\operatorname{PU}}(1,n)$ be a complex hyperbolic lattice and $\rho:\Gamma\to{\mathbf{H}}(k)$ be a rich representation. Then $\rho$ is complex (geodesically) rich if infinitely many of the $S_{\Gamma_{i}}$ for which $\dim{\mathbf{H}}_{i}<\dim{\mathbf{H}}$ are complex hyperbolic.
@@
Question 1.8.
Do rich representations extend? Equivalently, are there any nontrivial rich representations?
@@
Fix a parabolic subgroup $P\subset G$ with unipotent radical $U$. A pair $(k,{\mathbf{H}})$ as above is strongly compatible with $G$ if for any subgroup $\{1\}\neq{\mathbf{H}}^{\prime}\subset{\mathbf{H}}$ and every continuous group homomorphism $\tau:P\to N_{\mathbf{H}}({\mathbf{H}}^{\prime})/{\mathbf{H}}^{\prime}(k)$ we have that $U\leq\operatorname{Ker}(\tau)$.
@@
Our main theorem establishes 1.8 in two cases:
@@
    ${\mathbf{H}}$ is strongly compatible (in the sense of 1.9) and $k$-simple, or
@@
    $\Gamma$ is complex hyperbolic, $\rho$ is complex (geodesically) rich, and $\rho$ is cohomologically rigid (i.e., $H^{1}(\Gamma,\operatorname{Ad}_{{\mathbf{H}}}\circ\rho)=0$),
@@
Notice that the former condition is just on the pair $(k,{\mathbf{H}})$, whereas the latter depends only on $\rho$. The proof of (1), which follows from the methods developed in [1, 2], is in Section 2. The proof of (2), contained in Section 7, is related to the study of variations of Hodge structures and in fact it will be related to the so-called Zilber–Pink conjecture [7]. Regarding (2), it might be possible to study some nonrigid representations via the so called factorization theorem, see for example [37, Thm. 10] and various more recent generalizations. The fact that $(k,{\mathbf{H}})$ is always compatible when $k$ is nonarchimedean has the following fairly direct corollary.
@@
1.12 should be viewed as a complement to the fundamental recent work of Esnault and Groechenig [17] that proves integrality of certain cohomologically rigid local systems. Note that 1.12 does not assume cohomological rigidity, only that the representation is defined over a number field. Moreover, note that there are homomorphisms from complex hyperbolic lattices onto surface groups [40], hence not all representations of complex hyperbolic lattices are definable over a number field. As noticed in [2], the version of 1.12 that implicitly follows from the results there makes the main theorem of that paper independent of [17].
We close this subsection with a comment on how richness may differentiate between real and complex hyperbolic lattices. For example, we do not know the answer to the following:
Question 1.13.
@@
Bending deformations of lattices in $\operatorname{\operatorname{PO}}(1,n)$ [22, Sec. 5] easily imply that 1.13 is false for real hyperbolic lattices of simplest kind. However, Goldman–Millson rigidity [20] implies that there are no analogous deformations in the complex hyperbolic setting. It seems that examples giving a negative answer to 1.13 would come from representations of a new kind. On the other hand, a positive answer would (for example) prove superrigidity of unbounded, Zariski dense, and faithful representations to connected, adjoint targets.
Since 1.11 proves that all rich representations to connected, adjoint $k$-simple groups over nonarchimedean fields of characteristic zero are bounded, all other results in this paper consider the cases where $k$ is $\mathbb{R}$ or $\mathbb{C}$. For the remainder of this introduction we consider rich representations of complex hyperbolic lattices with coefficients in $\mathbb{R}$.

### v3

# Rich representations and superrigidity

@@ Abstract.
We investigate and compare applications of the Zilber–Pink conjecture and dynamical methods to rigidity problems for arithmetic real and complex hyperbolic lattices. Along the way we obtain new general results about reconstructing a variation of Hodge structure from its typical Hodge locus that may be of independent interest. Applications to Siu’s immersion problem are also discussed, the most general of which (1.6) only requires the hypothesis that infinitely many closed geodesics map to proper totally geodesic subvarieties under the immersion.
@@ 1. Introduction and Motivation
A cornerstone in the representation theory of lattices in Lie groups is the Margulis superrigidity theorem [31, p. 2]. Indeed, the implications of superrigidity are profound, including arithmeticity of irreducible lattices in all noncompact semisimple Lie groups except $\operatorname{\operatorname{PO}}(1,n)$ and $\operatorname{\operatorname{PU}}(1,n)$, i.e., for all but real and complex hyperbolic lattices. Specifically, Margulis proved superrigidity of irreducible lattices in semisimple Lie groups of real rank at least two. Using the theory of harmonic maps, this was later extended to the rank one groups $\mathrm{Sp}(1,n)$ and $\mathrm{F}_{4}^{(-20)}$ by Corlette [14] and Gromov–Schoen [22]. While the harmonic maps method was extended to give a new proof in higher rank [32], it is open (and quite relevant in the context of the blend of techniques used in this paper) whether the dynamical methods of Margulis can give an independent proof in the rank one cases.
To be precise, let $\Gamma$ be a lattice in a Lie group $G$. A representation $\rho$ from $\Gamma$ to a topological group $H$ is said to extend if there is a continuous homomorphism $\widetilde{\rho}:G\to H$ so that $\rho$ is the restriction of $\widetilde{\rho}$ to $\Gamma$ under its lattice embedding in $G$. Then $\Gamma$ is superrigid if every unbounded, Zariski dense representation of $\Gamma$ to a connected, adjoint simple algebraic group $H$ over a local field $k$ of characteristic zero extends. It is known that there are lattices in $\operatorname{\operatorname{PO}}(1,n)$ and $\operatorname{\operatorname{PU}}(1,n)$ that are not superrigid; [23, Sec. 5] for real hyperbolic examples and [34, Sec. 22] or [27, Sec. 1] for some results in the complex hyperbolic setting. The study of representations of real and complex hyperbolic lattices is very difficult and rich in open questions, and the extent to which superrigidity fails has yet to be made precise. For example, one motivation for this paper is the first part of a question asked by David Fisher [20, Qtn. 3.15]:
@@
In particular, this paper builds on the techniques from [1, 2, 8] to explore geometric conditions on a representation that allow one to use tools from dynamics and/or Hodge theory to prove new rigidity phenomena. The conditions we provide, which address 1.1 through the behavior of totally geodesic submanifolds under maps related to the representation, have also been considered in studying the infinite volume setting; see for example [26, 25]. There is some history of success along these lines. For just one family of examples, some classes of representations studied in the literature have maximal geometric behavior, where two distinct instances are:
@@
Since it arises several times, we briefly recall what it means for $\Gamma\subset G$ to be an arithmetic lattice of the simplest kind. These are well-studied in the literature, so we are terse. In the real hyperbolic case, this means that there is a totally real number field $F$ with integer ring $\mathcal{O}_{F}$ and a quadratic form $q$ on $F^{n+1}$ so that $\Gamma$ is commensurable with $\mathcal{G}(\mathcal{O}_{F})$, where $\mathcal{G}$ is the $F$-algebraic group $\operatorname{\operatorname{PO}}(q)$. In the complex hyperbolic case, there is a totally imaginary extension $E$ of the totally real field $F$ so that $\Gamma$ is commensurable with $\mathcal{G}(\mathcal{O}_{F})$, where $\mathcal{G}$ is now $\operatorname{\operatorname{PU}}(h)$ for a hermitian form $h$ on $E^{n+1}$ (which is still an $F$-algebraic group). In each case, it is necessarily the case that the form has signature $(1,n)$ at one place of $F$ and is definite at all other places. The associated locally symmetric spaces have an abundance of totally geodesic submanifolds, which are the main object of study in this paper.
@@ 1.1. Siu’s immersion problem
An original motivation for studying the more general rigidity questions considered in this paper is Siu’s immersion problem [41, Prob. (b), p. 182]. Let $\mathbb{B}^{n}$ denote the unit ball in $\mathbb{C}^{n}$ with its Bergman metric. If $\Gamma$ is a lattice in the holomorphic isometry group $\operatorname{\operatorname{PU}}(1,n)$ of $\mathbb{B}^{n}$, then $S_{\Gamma}$ will denote the ball quotient $\Gamma\backslash\mathbb{B}^{n}$.
@@
To our knowledge the only previous result is by Cao and Mok [13], which answered 1.2 in the affirmative for $f:\Gamma\backslash\mathbb{B}^{n}\to\Lambda\backslash\mathbb{B}^{m}$ when $m<2n$. See [2, Thm. 1.7] and the surrounding references for more on Siu’s analogous submersion problem. The following is our main contribution toward a positive answer to Siu’s immersion problem.
@@
Let $f:S_{\Gamma_{1}}\to S_{\Gamma_{2}}$ be an immersive holomorphic map between arithmetic ball quotients such that $f(S_{\Gamma_{1}})$ does not lie in any smaller totally geodesic subvariety of $S_{\Gamma_{2}}$. Suppose that $S_{\Gamma_{1}}$ contains a proper totally geodesic subvariety of positive dimension. Then the following are equivalent:
@@
The positive-dimensional hypothesis on the subvarieties of $S_{\Gamma_{1}}$ is to rule out the trivial counterexample of points. We offer two proofs of 1.3, one using homogeneous dynamics which gives a slightly weaker statement in Section 3, and the second using Hodge theory and functional transcendence in Section 4, as an application of 1.18 below.
For arithmetic lattices of simplest kind, 1.3 reduces 1.2 to checking a single dimension.
@@
There are various similar generalizations that follow from 1.17. For example, a period map $\psi:S_{\Gamma_{1}}\to\Gamma\backslash D$ such that the $f^{-1}(Y_{i})$ are totally geodesic subvarieties of $S_{\Gamma_{1}}$ for a generic sequence of Mumford–Tate subdomains $Y_{i}\subset\Gamma\backslash D$ is necessarily an isomorphism. Here ‘generic’ means that they are not all contained in a finite union of maximal totally geodesic subvarieties of $S$.
From a dynamical point of view, the dimension at least two hypothesis in the previous results is natural and critical in that it allows one to access tools from unipotent dynamics. From a Hodge theoretic point of view, one typically works with a complex subvariety. However, there is still something one can say for closed geodesics.
Theorem 1.6.
$f:S\longrightarrow\Gamma_{m}\backslash\mathbb{B}^{m}$
such that $f(S)$ does not lie in any smaller totally geodesic subvariety of $\Gamma_{m}\backslash\mathbb{B}^{m}$. Let $\{C_{\ell}\}$ be a generic sequence of distinct closed geodesics in $S$ in the sense that they are not all contained in a finite union of maximal totally geodesic subvarieties of $S$. Assume that for each $\ell$ there exits a strict complex totally geodesic subvariety $Y_{\ell}\subset\Gamma_{m}\backslash\mathbb{B}^{m}$ such that $f(C_{\ell})\subset Y_{\ell}$. Then $f$ is a totally geodesic immersion.
See 4.4 for the proof. Note that 1.6 applies for nonarithmetic $\Gamma$, since all finite volume ball quotients contain infinitely many free homotopy classes of closed geodesics. That $f$ must take certain closed geodesics into a proper totally geodesic subspace is a very strong hypothesis, but it is not entirely unprecedented. Indeed one can show that some of the maps associated with surjections between complex hyperbolic lattices take a certain (necessarily finite) collection of totally geodesic subspaces of $S_{\Gamma_{n}}$ to totally geodesic subspaces of $S_{\Gamma_{m}}$. We close this section with a final question that would generalize 1.6.
Question 1.7.
@@ 1.2. Rich representations and the main theorem
Let $\Gamma\subset G$ be a torsion-free arithmetic lattice, where $G$ is $\operatorname{\operatorname{SU}}(1,n)$ or $\operatorname{\operatorname{SO}}(1,n)$, and let $X$ denote the symmetric space associated with $G$. Then $S_{\Gamma}$ will denote the manifold $\Gamma\backslash X$, which is a smooth manifold that is also a smooth variety in the complex hyperbolic case [^1] Since both contexts are relevant to this paper, we use the term submanifold when considering $S_{\Gamma}$ as a differentiable manifold and subvariety when $S_{\Gamma}$ is complex hyperbolic and the submanifold is also a subvariety for the natural algebraic structure.. We assume that $S_{\Gamma}$ contains one, and therefore infinitely many, proper totally geodesic submanifolds [^2] This claim follows easily from the fact that the associated algebraic group commensurates the lattice. In the other direction, if $S_{\Gamma}$ contains infinitely many maximal totally geodesic submanifolds of dimension at least $2$, then $\Gamma$ is arithmetic; see [1, 2, 8]. of real dimension at least $2$. When this is the case, we say that $S_{\Gamma}$ (or $\Gamma$) is geodesically rich. Let $\{S_{\Gamma_{i}}\}_{i\in\mathbb{N}}$ be a sequence of distinct immersed maximal totally geodesic submanifolds, where maximal means that the only totally geodesic submanifold strictly containing $S_{\Gamma_{i}}$ is $S_{\Gamma}$ itself.
The group $\Gamma_{i}$ is the stabilizer in $\Gamma$ of a totally geodesic subspace $Y_{i}$ of $X$, and the restriction of $\Gamma_{i}$ to $Y_{i}$ induces a surjective homomorphism from $\Gamma_{i}$ onto the fundamental group of $S_{\Gamma_{i}}$. The kernel of this homomorphism encodes the action of $\Gamma_{i}$ on the normal bundle to $Y_{i}$ in $X$. The main purpose of this paper is to introduce the notion of geodesically rich representations, which are morally the ones that “preserve” a significant amount of the rich structure of $\Gamma$:
@@
Let $k$ be a local field and ${\mathbf{H}}$ be a semisimple $k$-algebraic group without compact factors. Let $\rho:\Gamma\to{\mathbf{H}}(k)$ be a representation with unbounded, Zariski dense image. We say that $\rho$ is geodesically rich (just rich when it is clear from context) if the following are satisfied:
  • $\Gamma$ is geodesically rich, and
  • there are infinitely many $i$ so that $\dim{\mathbf{H}}_{i}<\dim{\mathbf{H}}$, where ${\mathbf{H}}_{i}$ is defined by
    (1.1) $\mathbf{H}_{i}\coloneqq\overline{\rho(\Gamma_{i})}\subset{\mathbf{H}}$
    for each $i$, where closure is with respect to the Zariski topology.
@@
Let $\Gamma\subset G=\operatorname{\operatorname{PU}}(1,n)$ be a complex hyperbolic lattice and $\rho:\Gamma\to{\mathbf{H}}(k)$ be a rich representation. Then $\rho$ is complex geodesically rich if infinitely many of the $S_{\Gamma_{i}}$ for which $\dim{\mathbf{H}}_{i}<\dim{\mathbf{H}}$ are complex hyperbolic.
@@
Question 1.10.
Do rich representations extend? Equivalently, are there nontrivial rich representations (i.e., other than the case where ${\mathbf{H}}(k)=G$ and $\rho$ is given by conjugation)?
@@
Fix a parabolic subgroup $P\subset G$ with unipotent radical $U$. A pair $(k,{\mathbf{H}})$ as above is strongly compatible with $G$ if for any subgroup $\{1\}\neq{\mathbf{H}}^{\prime}\subset{\mathbf{H}}$ and every continuous group homomorphism $\tau:P\to(N_{\mathbf{H}}({\mathbf{H}}^{\prime})/{\mathbf{H}}^{\prime})(k)$ we have that $U\leq\operatorname{Ker}(\tau)$.
@@
Our main theorem establishes 1.10 in two cases:
@@
    ${\mathbf{H}}$ is strongly compatible (in the sense of 1.11) and $k$-simple, or
@@
    $\Gamma$ is complex hyperbolic, $\rho$ is complex geodesically rich, and $\rho$ is cohomologically rigid (i.e., $H^{1}(\Gamma,\operatorname{Ad}_{{\mathbf{H}}}\circ\rho)=0$),
@@
Notice that the former condition is just on the pair $(k,{\mathbf{H}})$, whereas the latter depends only on $\rho$. The proof of (1), which follows from the methods developed in [1, 2], is in Section 2. The proof of (2), contained in Section 7, is related to the study of variations of Hodge structures and in fact it will be related to the so-called Zilber–Pink conjecture [7]. Regarding (2), it might be possible to study some nonrigid representations via the so called factorization theorem, see for example [40, Thm. 10] and various more recent generalizations. The fact that $(k,{\mathbf{H}})$ is always compatible when $k$ is nonarchimedean has the following fairly direct corollary.
@@
1.14 should be viewed as a complement to the fundamental recent work of Esnault and Groechenig [18] that proves integrality of certain cohomologically rigid local systems. Note that 1.14 does not assume cohomological rigidity, only that the representation is defined over a number field. Moreover, note that there are homomorphisms from complex hyperbolic lattices onto surface groups [43], hence not all representations of complex hyperbolic lattices are definable over a number field. As noticed in [2], the version of 1.14 that implicitly follows from the results there makes the main theorem of that paper independent of [18].
Remark 1.15.
In contrast with 1.6, we now give a simple cautionary example regarding generalizing richness to closed geodesics. Let $G$ be $\operatorname{\operatorname{PO}}(1,n)$ for $n\geq 3$ or $\operatorname{\operatorname{PU}}(1,n)$ for $n\geq 2$, and suppose that $\Gamma_{1},\Gamma_{2}$ are arithmetic lattices in $G$ of simplest type for which there is a surjection $\rho:\Gamma_{1}\to\Gamma_{2}$ with infinite kernel; such representations are known to exist in both settings. Choose a maximal subgroup $\Delta\subset\Gamma_{2}$ associated with a proper totally geodesic submanifold of $S_{\Gamma_{2}}$. Then $\rho^{-1}(\Delta)$ contains the kernel of $\rho$, and therefore is Zariski dense in $G$. Consequently, $\rho^{-1}(\Delta)$ contains a profinitely open collection of regular semisimple elements [37], hence they preserve no totally geodesic subspace of $X$ except their axis of translation, and thus the associated closed geodesics on $S_{\Gamma}$ are maximal. In the complex hyperbolic setting the associated closed geodesics are generic in the sense that their union is Zariski dense in $S_{\Gamma}$, hence this provides a fairly rich collection of $1$-dimensional maximal totally geodesic subspaces.
Representations as in 1.15 are unbounded and Zariski dense with connected, adjoint target, but they cannot extend since they have infinite kernel. They have a well-distributed sequence of maximal closed geodesics for which the associated elements in $\rho(\Gamma_{1})$ all preserve a totally geodesic subspace of the target, hence the representation could be considered rich in analogy with the higher-dimensional setting. While one might reasonably interpret this as saying the $1$-dimensional case is hopeless, 1.6 indicates that stronger geometric hypotheses can still lead to a rigidity theorem.
We now comment on how richness may differentiate between real and complex hyperbolic lattices. For example, we do not know the answer to the following:
Question 1.16.
@@
Bending deformations of lattices in $\operatorname{\operatorname{PO}}(1,n)$ [23, Sec. 5] easily imply that 1.16 is false for real hyperbolic lattices of simplest kind. However, Goldman–Millson rigidity [21] implies that there are no analogous deformations in the complex hyperbolic setting. It seems that examples giving a negative answer to 1.16 would come from representations of a new kind. On the other hand, a positive answer would (for example) prove superrigidity of unbounded, Zariski dense, and faithful representations to connected, adjoint targets.
Since 1.13 proves that all rich representations to connected, adjoint $k$-simple groups over nonarchimedean fields of characteristic zero are bounded, all other results in this paper consider the cases where $k$ is $\mathbb{R}$ or $\mathbb{C}$. For the remainder of this introduction we consider rich representations of complex hyperbolic lattices with coefficients in $\mathbb{R}$.